OpenStudy (anonymous):

Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar? I need some help with this I"m really confused really I koe formulas i just need to know HOW to solve

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

match volumes

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

n*volume of 1 penny = volume of jar... solve for n (number of pennies which should theoretically fit in the jar).

OpenStudy (anonymous):

n*0.375*0.375*3.14*0.061=3*3*3.14*11.5?

OpenStudy (anonymous):

\[n =\frac{\text{volume of jar}}{\text{volume of a penny}}\]

OpenStudy (anonymous):

\[n=\frac{ \cancel{\pi}\cdot 3^2 \cdot 11.5 }{ \cancel{\pi} \cdot \left( 0.375 \right)^2 \cdot 0.061 }\]

OpenStudy (anonymous):

thats such a big number

OpenStudy (anonymous):

that's a bunch of pennies

OpenStudy (anonymous):

well thas abount 12 thousand with many other pennys

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

wel thats not realistic are you sure?

OpenStudy (anonymous):

sure it is. while you're ultimately not going to be able to fill the entire volume with pennies (there will be air gaps) you could fit alot of pennies in a jar that size. it's nearly a foot tall and half a foot in diameter!

OpenStudy (anonymous):

3.14*3*3*11.5/3.14*0.375*0.375*0.061=324.99/0.0269353125=12065.573770491803278688524590164=12065? corect?

OpenStudy (anonymous):

yep, that's what i got

OpenStudy (anonymous):

ok